# latexfontFont shape Assumptions of Randomness in Cosmology Models

### v3

@@ Abstract
Non-compact symmetries cannot be fully broken by randomness since
non-compact
@@
Physics laws allow all sorts of events and histories. In infinite cosmology models each one of them rolls out infinitely often: weirdest miracles are daily routine in some places. To understand our boring world one must assume our place is rather generic. So, theories assume some events happen at random, others are derived from them by physical equations or other algorithms. In fact, [Gács 86, Kučera 85] prove that any sequence is computable from an algorithmically random one.111Algorithmic randomness (see: [Kolmogorov 65, Kolmogorov, Uspenskii 87]) of a digital sequence does not quite assure it obeys all probability math laws (only computable ones). Yet, it assures (see: [Levin 84]) that either obeys all such laws or has an infinite information about some uncomputable but mathematically definable object. The latter case would be really weird. There are no ways for such to be generated; see this informational version of Church-Turing Thesis discussed at the end of [Levin 13]. However, the theories of this type can only be complete if they specify which exactly are the random events and what is their probability distribution. These events need not be directly observable, rather just be a part of the model from which the observations are derived. By [V’yugin 82], resulting observables, unlike random ones, could be very diverse in algorithmically invariant properties.
@@
A technical difficulty comes up here. While, the observable part of the world is finite, many models include it in a greater infinite realm, beyond our reach. How to choose our (observer’s) place in this realm at random is puzzling. Say, in a most primitive setting, how to choose a random point in a pattern on an infinite Euclidean space? It carries no uniform probability distribution. (In other models, where ours is one of the infinitely many “bubbles,” it may be even much trickier.) A way around this technicality can lie in the reverse order of choices. Instead of first building an infinite universe and then choosing the place of the observer there, we can first designate the observer, then build an infinite universe around her.333One may argue that computing probability of any finite event requires only placing the observer in a finite region, which can be done at random. However this does not provide a complete model of the infinite universe. Some global properties do not manifest in any finite region. And a compact event is the intersection of a monotone family of clopen (finite) events . Its probability is Prob. But s must all agree on the observer’s surroundings. So that would anyway amount to setting the observer first and then using randomness to generate the universe around. This circumvents the lack of uniform distributions in infinite domains (and tickles our ego, besides :-). Such structures are called pointed spaces in topology. Among the challenges is the need to expand the universe not only to the observer’s present and all future times, but also to her whole past (if there are any causal relations with the past).
@@
This brings up a question, what is an observer? Its many meanings depend on what symmetries are meant to be broken by choosing this sort of a “gauge”. In quantum physics observer can mean a macroscopic system brought into an entanglement with the observed microscopic variables. “Macroscopic” is a vague term with unclear relevance. Sometimes it is put in relation with the Plank mass (e.g., by R.Penrose). But it is unclear why a microscopic bacteria, much lighter than Plank mass, cannot perfectly serve as an observer. (It could transmit its observed data to its descendants, which may include members of Royal Society. :-)
@@
What seems relevant here, is that the observed data, unlike generic quantum states, can be reliably copied, preserved, transmitted, etc. For this an observer needs some sophistication, access to mechanisms for error-correction, self-preservation, etc. (This requires free energy flow: so gravity, its ultimate source, may play a role for observers, even if not via Plank mass. To enter a mixed state, observers need to discard parts of the system. Thus they consume free energy and expel heat to some horizon: cosmological or black hole’s, or some such. This generates entropy, both of the observer and of the horizon if taken separately, though commonly used entropy formulas for those horizons are somewhat speculative.444When a max mass neutron star absorbs a little more gas and collapses into a black hole, its mass changes little, radius shrinks just twice or so, but its entropy grows by 20 orders of magnitude, approaching the combined entropy of all neutron stars in the universe. This sounds strange. Ordinary matter systems cannot have entropy in volume less than about in Planck units: by Pauli exclusion for fermions, Stefan-Boltzmann energy density for photons, Schwarzschild radius limit, etc. But black holes of such volume get Bekenstein-Hawking entropy (= area), the bound inspired by a GR theorem that the total black holes surface areas cannot shrink. I do not know if GR has counterexamples to a stronger statement that the sum of square roots of black holes volumes never shrinks either, or if LIGO can find colliding black holes that radiate away over 20% of their combined mass.) Such features are readily present in life.
@@
fully Anthropic? :-) I assume for observers only the seemingly

### v4

# Assumptions of Randomness in Cosmology Models

@@ Abstract
Non-compact symmetries cannot be fully broken by randomness
since non-compact
@@
Physics laws allow all sorts of events and histories. In infinite cosmology models each one of them rolls out infinitely often: weirdest miracles are daily routine in some places. To understand our boring world one must assume our place is rather generic. So, theories assume some events happen at random, others are derived from them by physical equations or other algorithms. In fact, [Gács 86, Kučera 85] prove that any sequence is computable from an algorithmically random one.111Algorithmic randomness (see: [Kolmogorov 65, Kolmogorov, Uspenskii 87]) of a digital sequence does not quite assure it obeys all probability math laws (only computable ones). Yet, it assures (see: [Levin 84]) that either obeys all such laws or has an infinite information about some mathematically definable object. The latter case would be really weird. There are no ways for such to be generated; see this informational version of Church-Turing Thesis discussed at the end of [Levin 13]. However, the theories of this type can only be complete if they specify which exactly are the random events and what is their probability distribution. These events need not be directly observable, rather just be a part of the model from which the observations are derived. By [V’yugin 82], resulting observables, unlike random ones, could be very diverse in algorithmically invariant properties.
@@
A technical difficulty comes up here. While the observable part of the world is finite, many models include it in a greater infinite realm, beyond our reach. How to choose our (observer’s) place in this realm at random is puzzling. Say, in a most primitive setting, how to choose a random point in a pattern on an infinite Euclidean space? It carries no uniform probability distribution. (In other models, where ours is one of the infinitely many “bubbles,” it may be even much trickier.) A way around this technicality can lie in the reverse order of choices. Instead of first building an infinite universe and then choosing the place of the observer there, we can first designate the observer, then build an infinite universe around her.333One may argue that computing probability of any finite event requires only placing the observer in a finite region, which can be done at random. However this does not provide a complete model of the infinite universe. Some global properties do not manifest in any finite region. And a compact event is the intersection of a monotone family of clopen (finite) events . Its probability is Prob. But s must all agree on the observer’s surroundings. So that would anyway amount to setting the observer first and then using randomness to generate the universe around. This circumvents the lack of uniform distributions in infinite domains (and tickles our ego, besides :-). Such structures are called pointed spaces in topology. Among the challenges is the need to expand the universe not only to the observer’s present and all future times, but also to her whole past (if there are any causal relations with the past).
@@
This brings up a question, what is an observer? Its many meanings depend on what symmetries are meant to be broken by choosing this sort of a “gauge”. In quantum physics observer can mean a macroscopic system brought into an entanglement with the observed microscopic variables. “Macroscopic” is a vague term with unclear relevance. Sometimes it is put in relation with the Planck mass (e.g., by R.Penrose). But it is unclear why a microscopic bacteria, much lighter than Planck mass, cannot perfectly serve as an observer. (It could transmit its observed data to its descendants, which may include members of Royal Society. :-)
@@
What seems relevant here, is that the observed data, unlike generic quantum states, can be reliably copied, preserved, transmitted, etc. For this an observer needs some sophistication, access to mechanisms for error-correction, self-preservation, etc. (This requires free energy flow: so gravity, its ultimate source, may play a role for observers, even if not via Planck mass. To enter a mixed state, observers need to discard parts of the system. Thus they consume free energy and expel heat to some horizon: cosmological or black hole’s, or some such. This generates entropy, both of the observer and of the horizon if taken separately, though commonly used entropy formulas for those horizons are somewhat speculative.444When a max mass neutron star absorbs a little more gas and collapses into a black hole, its mass changes little, radius shrinks just twice or so, but its entropy grows by 20 orders of magnitude, approaching the combined entropy of all neutron stars in the universe. This sounds strange. Ordinary matter systems cannot have entropy in volume less than about in Planck units: by Pauli exclusion for fermions, Stefan-Boltzmann energy density for photons, Schwarzschild radius limit, etc. But black holes of such volume get Bekenstein-Hawking entropy (= area), the bound inspired by a GR theorem that the total black holes surface areas cannot shrink. I do not know if GR has counterexamples to a stronger statement that the sum of square roots of black holes volumes never shrinks either, or if LIGO can find colliding black holes that radiate away over 20% of their combined mass.) Such features are readily present in life.
@@
fully Anthropic? (:-) I assume for observers only the seemingly